# NBBOX: Noisy Bounding Box Improves Remote Sensing Object Detection

### v1

@@ Abstract
Data augmentation has seen significant advancements in computer vision to improve model performance over the years, particularly in scenarios with limited and insufficient data. Currently, most studies focus on adjusting the image or its features to expand the size, quality, and variety of samples during training in various tasks including object detection. However, we argue that it is necessary to investigate bounding box transformations as a model regularization technique rather than image-level transformations, especially in aerial imagery due to potentially inconsistent bounding box annotations. Hence, this letter presents a thorough investigation of bounding box transformation in terms of scaling, rotation, and translation for remote sensing object detection. We call this augmentation strategy NBBOX (Noise Injection into Bounding Box). We conduct extensive experiments on DOTA and DIOR-R, both well-known datasets that include a variety of rotated generic objects in aerial images. Experimental results show that our approach significantly improves remote sensing object detection without whistles and bells and it is more time-efficient than other state-of-the-art augmentation strategies.
@@ I Introduction
In recent years, remote sensing object detection has achieved remarkable success primarily due to the advancements in deep learning. Particularly, modern deep learning architectures such as convolutional neural networks and transformers have significantly enhanced the capabilities of object detection in Earth vision. These models have millions or billions of parameters to learn, requiring tremendous training data to avoid the over-fitting problem. Compared to natural scenes (e.g. ImageNet [1] or MS-COCO [2]), datasets for overhead imagery usually lack diversity and quantity due to the expense of data collection. To overcome this limitation, data augmentation plays a significant role as an implicit regularization strategy for model learning.
@@
The main contributions of this work are as follows:
  • To the best of our knowledge, we first investigate the bounding box transformation to improve remote sensing object detection. With our investigation, we provide a useful recipe for adding noise to bounding boxes.
  • It is worth noting that our method is easy to plug in and merely increases the computational complexity compared to other state-of-the-art data augmentation strategies.
  • Extensive experimental results demonstrate that NBBOX can boost remote sensing object detection without whistles and bells, against implicit box label noise.
@@ II Methodology
In this section, we present a simple, but efficient data augmentation strategy namely NBBOX (Noise Injection into Bounding Box) for remote sensing object detection. Our method can be easily integrated into existing detection frameworks with minimal modifications. More precisely, our augmentation method can be seamlessly added to any data pre-processing pipeline without severe computational burden. The proposed approach is straightforward as depicted in Fig. 1. For data augmentation, our training scheme randomly adds noise to oriented bounding boxes with the three simplest geometric transformations: scaling, rotation, and translation.
@@ II-A Background
For remote sensing images, a few work has been proposed to address noisy labels in object detection, inspired by research for natural scenes. As seen in Fig. 2, there is a discrepancy between the actual annotation and the minimum enclosing box, which might cause difficulties in model training due to inaccurate and low-consistent bounding boxes. [19] designs two kinds of loss functions to mitigate mislabels for both classification and localization. On the other hand, [20], [21] focus on correcting incomplete annotations such as missing bounding boxes due to misalignment of GPS sensors. The main goal of our work is to resolve localization errors inherent in the existing bounding box labels. In other words, we consider a generic object detection scenario, where missing bounding boxes or incorrect class labels of the ground truth are uncommon. Instead, we assume that the bounding box labels may not be optimal due to the annotators’ mistakes or dilemma situations such as how to rotate a bounding box for a circular object on overhead imagery during the annotation process. In particular, we study a data augmentation strategy based on bounding box transformation for more robust localization.
@@ II-C Advanced options for translation and scaling
As demonstrated in Algorithm 1, the bounding box may be translated by different values along both the x-axis and y-axis. Besides, the aspect ratio may not be maintained for scaling operations. This phenomenon would not be recommended for training on certain data scenarios. For instance, it would not be desirable to modify the aspect ratio of bounding boxes in an extreme situation where most objects are expected to have a square size. Hence, we alleviate this by providing additional parameters ${\text{bool}}_{s}\in\{\text{True},\text{False}\}$ and ${\text{bool}}_{t}\in\{\text{True},\text{False}\}$ in our NBBOX algorithm. Here, ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ are boolean values to be set by users. One can effortlessly modify our algorithm with ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ as follows.
@@ III-A Experimental setup
Detection models. Traditionally, deep learning-based object detection models are typically separated into one-stage and two-stage methods. Nowadays, the distinction between anchor-based and anchor-free methodologies tends to be more emphasized when classifying detector models. In this work, we adopt Faster R-CNN [22], RetinaNet [23], and FCOS [24] which are representative models in anchor-based two-stage, anchor-based one-stage, and anchor-free one-stage detection studies, respectively. For all models, Imagenet-pretrained ResNet-50 [25] is used for feature extraction, and FPN-1x [26] is added as a neck to improve multi-scale invariance.
@@
Training details. We implement our hypothesis using PyTorch and MMRotate [27]. All the models are trained for 25 epochs. For a fair comparison with other approaches, all the experiments are resumed with weights pre-trained for five epochs without any augmentation per each {detector, data} combination. SGD is used for optimization with momentum of weight 0.9, weight decay 1e-4, learning rate 1e-3, and batch size 8. Learning rate warm-up is used over the first 500 iterations, where the learning rate linearly increases from zero to 33% of 5e-3. L2-norm gradient clipping is utilized with a maximum norm of 35. We apply the following pre-processing operations for training: ‘RResize’, ‘RRandomFlip’, ‘Normalize’, and ‘Pad’, while for testing, ‘MultiScaleFlipAug’ is also used. For the sake of brevity and due to time limitations, we only conduct the parameter search on DIOR-R and scale the best configurations to DOTA in this work.
@@
Figure 3: Impact of scaling, rotation, and translation of bounding boxes on model performance: as we observe similar outcomes across different architectures and datasets, for brevity, we include only the result from Faster R-CNN trained on DIOR-R.
TABLE I: Experimental results of scaling, rotation, and translation of bounding boxes across various detection architectures on DIOR-R with NBBOX.
(Bold text means the best while underlined and wavy-underlined text indicate the second- and third-best, respectively in this work)
Configuration Detectors
Transform
${s}_{\text{min}}$
(${r}_{\text{min}}$)
<${t}_{\text{min}}$>
${s}_{\text{max}}$
(${r}_{\text{max}}$)
<${t}_{\text{max}}$>
${\text{bool}}_{s}$
<${\text{bool}}_{t}$>
Faster R-CNN FCOS RetinaNet
Scaling 0.5x 1x True 51.52 ± 0.34 42.17 ± 1.04 36.92 ± 0.36
False 51.52 ± 0.34 43.41 ± 1.17 37.90 ± 0.88
0.8x 1x True 60.49 ± 0.21 58.50 ± 0.88 54.48 ± 0.60
False 60.52 ± 0.45 58.52 ± 1.03 54.10 ± 1.02
0.9x 1x True 62.36 ± 0.28 59.60 ± 1.01 55.46 ± 0.72
False 62.39 ± 0.22 59.39 ± 0.76 55.06 ± 0.68
0.95x 1x True 62.66 ± 0.20 59.78 ± 0.78 55.60 ± 0.65
False 62.45 ± 0.11 59.58 ± 0.92 55.77 ± 0.82
0.99x 1x True 62.79 ± 0.05 60.00 ± 0.99 55.50 ± 0.55
False 62.65 ± 0.18 59.74 ± 1.13 55.31 ± 0.85
1x 1.01x True 62.57 ± 0.14 59.91 ± 0.86 55.79 ± 0.72
False 62.71 ± 0.27 59.68 ± 1.09 55.74 ± 0.47
1x 1.05x True 62.63 ± 0.32 59.68 ± 0.99 55.64 ± 0.75
False 62.28 ± 0.09 59.55 ± 1.12 55.40 ± 0.78
1x 1.1x True 62.41 ± 0.21 59.56 ± 1.15 55.33 ± 0.70
False 62.12 ± 0.21 59.48 ± 0.97 55.44 ± 0.62
1x 1.2x True 61.12 ± 0.24 58.77 ± 0.99 54.64 ± 0.69
False 61.03 ± 0.22 58.58 ± 0.95 54.60 ± 0.75
1x 1.5x True 52.68 ± 0.32 51.72 ± 1.14 48.38 ± 0.84
False 53.12 ± 0.21 51.28 ± 0.90 48.17 ± 0.80
(Rotation) -0.01 0.01 - 62.87 ± 0.29 59.68 ± 1.03 55.40 ± 0.53
-0.1 0.1 - 62.66 ± 0.15 59.66 ± 1.03 55.37 ± 0.90
-0.5 0.5 - 61.35 ± 0.21 58.46 ± 0.91 54.40 ± 0.99
-1 1 - 57.60 ± 0.26 54.16 ± 0.73 52.23 ± 0.41
-2 2 - 51.52 ± 0.34 37.95 ± 3.70 32.72 ± 2.23
-3 3 - 51.52 ± 0.34 39.04 ± 2.08 37.59 ± 0.53
-4 4 - 54.19 ± 0.27 47.85 ± 0.42 51.79 ± 0.85
-5 5 - 51.52 ± 0.34 38.88 ± 2.26 37.94 ± 0.42
-10 10 - 51.52 ± 0.34 43.44 ± 0.44 48.73 ± 0.23
<Translation> -1px 1px True 62.56 ± 0.12 59.93 ± 0.91 55.71 ± 0.74
False 62.73 ± 0.25 59.70 ± 0.95 55.42 ± 0.58
-2px 2px True 62.78 ± 0.26 59.72 ± 0.92 55.33 ± 0.69
False 62.59 ± 0.25 59.75 ± 1.02 55.45 ± 0.90
-3px 3px True 62.30 ± 0.19 59.62 ± 1.19 55.51 ± 0.61
False 62.56 ± 0.14 59.71 ± 0.88 55.69 ± 0.60
-4px 4px True 62.03 ± 0.20 59.46 ± 1.27 55.35 ± 0.70
False 61.73 ± 0.22 59.14 ± 1.39 55.48 ± 0.66
-5px 5px True 61.51 ± 0.28 59.49 ± 1.31 55.34 ± 0.74
False 61.78 ± 0.11 58.82 ± 1.41 55.39 ± 0.69
-10px 10px True 60.04 ± 0.15 58.33 ± 1.30 54.36 ± 1.08
False 59.23 ± 0.27 57.31 ± 1.37 53.99 ± 0.73
-20px 20px True 56.53 ± 0.22 56.15 ± 1.09 51.22 ± 1.05
False 52.17 ± 0.16 52.14 ± 0.99 48.88 ± 0.76
-30px 30px True 55.23 ± 0.36 54.30 ± 1.34 49.96 ± 0.64
False 51.52 ± 0.34 48.77 ± 0.88 45.81 ± 0.81
@@ III-C Impact of scaling, rotation, and translation in NBBOX
This subsection studies the model performance when exposed to varying levels of scaling, rotation, and translation of bounding boxes. For this, we consider combinations of the following parameters: ${s}_{\text{min}}\in$0.5, 0.8, 0.9, 0.95, 0.99}, ${s}_{\text{max}}\in$1, 1.01, 1.05, 1.1, 1.2, 1.5}, ${\text{bool}}_{s}\in$True, False}, $\{{r}_{\text{min}},{r}_{\text{max}}\}\in$±0.01, ±0.1, ±0.5, ±1, ±2, ±3, ±4, ±5, ±10}, and $\{{t}_{\text{min}},{t}_{\text{max}}\}\in$±1, ±2, ±3, ±4, ±5, ±10, ±20, ±30}, ${\text{bool}}_{t}\in$True, False}. As indicated in Fig. 3 and Table I, following hyper-parameters are recommended:
  • ${s}_{\text{min}}\in\{0.95,1\}\wedge{s}_{\text{max}}\in\{1,1.01\}$;
  • $\{{r}_{\text{min}},{r}_{\text{max}}\}=\{\pm 0.01\}$ ${t}_{\text{min}}=\pm 1$, ${t}_{\text{max}}=\pm 2$.
For rotation, we need to use fairly small parameter values, presumably as rotating bounding boxes can easily harm foreground content in aerial imagery, especially for tiny objects.
Isotropic scaling and translation. When scaling with preserving aspect ratio (i.e. ${\text{bool}}_{s}$=True), all models show a slight increase in mAP compared to when it is not (i.e. ${\text{bool}}_{s}$=False) as shown in Table II. Besides, in the case of translation noise, the models perform slightly better when isotropic translation is applied (i.e. ${\text{bool}}_{t}$=True) compared to when it is not (i.e. ${\text{bool}}_{t}$=False) as indicated in Table II. Thus, experimental results on DIOR-R in Table II suggest that isotropic scaling and translation may be important. However, it is vital to note that the performance observed in this experiment might not be consistently replicable across different datasets.
TABLE II: Average mAP under isotropic/non-isotropic scaling and
@@
Configuration Faster RCNN FCOS RetinaNet
Scaling ${\text{bool}}_{s}$=True 60.12 (+0.05) 56.97 (+0.05) 52.77 (+0.03)
${\text{bool}}_{s}$=False 60.07 (+0.00) 56.92 (+0.00) 52.74 (+0.00)
Translation ${\text{bool}}_{t}$=True 60.37 (+1.08) 58.38 (+1.46) 54.10 (+0.83)
${\text{bool}}_{t}$=False 59.29 (+0.00) 56.92 (+0.00) 53.27 (+0.00)
Combining scaling, rotation, and translation. There is a reasonable expectation that integrating scaling, rotation, and translation could synergistically improve the model’s robustness and generalization. Hence, we further conduct experiments to validate the effectiveness of such combinations. Precisely, we study different combinations of each transformation (with parameters that show the best): pairwise combinations and the complete combination. Table III indicates the effectiveness of the combined transformations in enhancing the model’s performance. Notably, the improvements are particularly evident in the scenario when we simultaneously use scaling, rotation, and translation during training.
@@
TABLE III: Impact of combining scaling, rotation, and translation of bounding boxes on DIOR-R in NBBOX.
Transform Faster R-CNN FCOS RetinaNet
Scaling (Best) 62.79 ± 0.05 60.00 ± 0.99 55.79 ± 0.72
Rotation (Best) 62.87 ± 0.29 59.68 ± 1.03 55.40 ± 0.53
Translation (Best) 62.78 ± 0.26 59.93 ± 0.91 55.71 ± 0.74
Scaling + Rotation 62.92 ± 0.24 60.13 ± 0.97 56.02 ± 0.65
Scaling + Translation 62.89 ± 0.30 60.02 ± 1.07 56.06 ± 0.65
Translation + Rotation 62.80 ± 0.22 59.89 ± 1.06 56.00 ± 0.72
Scaling + Rotation + Translation 63.18 ± 0.14 60.25 ± 1.00 56.09 ± 0.61
III-D Comparison with state-of-the-art methods
In this subsection, we compare our method with easy-to-get data augmentation methods based on image manipulation (image/patch erasing and mix) as illustrated in Table IV. We also compare ours with RandRotate and RandShift to show the difference (Our method differs as it only transforms bounding boxes while RandRotate and RandShift rotate or shift both the image and the bounding boxes). For a DOTA dataset, we follow the best parameters across three detectors on DIOR-R by hard voting. Additionally, we train and test only on Faster R-CNN, the best architecture on DIOR-R in our previous experiments. Table IV shows that our NBBOX achieves the best performance among the state-of-the-art augmentation methods. Though our method, when applied solely to DOTA, achieved competitive performance, ranking second behind RandRotate, its performance significantly improves when integrated with RandRotate. Such is true of RetinaNet on DIOR-R with MosaicMix. Besides, our method is faster than other image manipulation methods. Table V indicates the time required per each training epoch on Faster R-CNN for each augmentation. (e.g. for one epoch, ours vs MosaicMix: 9.63min vs 21.24min on DIOR-R). Due to the page limit, detailed settings are described in the supplementary material.

### v3

@@ Abstract
Data augmentation has shown significant advancements in computer vision to improve model performance over the years, particularly in scenarios with limited and insufficient data. Currently, most studies focus on adjusting the image or its features to expand the size, quality, and variety of samples during training in various tasks including object detection. However, we argue that it is necessary to investigate bounding box transformations as a data augmentation technique rather than image-level transformations, especially in aerial imagery due to potentially inconsistent bounding box annotations. Hence, this letter presents a thorough investigation of bounding box transformation in terms of scaling, rotation, and translation for remote sensing object detection. We call this augmentation strategy NBBOX (Noise Injection into Bounding Box). We conduct extensive experiments on DOTA and DIOR-R, both well-known datasets that include a variety of rotated generic objects in aerial images. Experimental results show that our approach significantly improves remote sensing object detection without whistles and bells and it is more time-efficient than other state-of-the-art augmentation strategies.
@@ I Introduction
In recent years, remote sensing object detection has achieved remarkable success primarily due to the advancements in deep learning. Particularly, modern deep learning architectures such as convolutional neural networks and transformers have significantly enhanced the capabilities of object detection in Earth vision (e.g. [1, 2]). These models have millions or billions of parameters to learn, requiring tremendous training data to avoid the over-fitting problem. Compared to natural scenes (e.g. ImageNet [3] or MS-COCO [4]), datasets for overhead imagery usually lack diversity and quantity due to the expense of data collection. To overcome this limitation, data augmentation plays a significant role as an implicit regularization strategy for model learning.
@@ II Methodology
In this section, we present a simple, but efficient data augmentation strategy namely NBBOX (Noise Injection into Bounding Box) for remote sensing object detection. As seen in Fig. 2, there is often a discrepancy between the actual annotation and the minimum enclosing box (i.e. optimal localization label) on the existing remote sensing datasets, which might cause difficulties in model training due to inaccurate and low-consistent bounding boxes. To enhance the model’s robustness against such implicit box-level noise, our augmentation scheme randomly adds noise to oriented bounding boxes with the three simplest geometric transformations: scaling, rotation, and translation during training, as depicted in Fig. 1.
@@ II-A Background
For remote sensing images, a few work has been proposed to address noisy labels in object detection, inspired by research for natural scenes. [21] designs two kinds of loss functions to mitigate mislabels for both classification and localization. On the other hand, [22], [23] focus on correcting incomplete annotations such as missing bounding boxes due to misalignment of GPS sensors. Unlike prior work, the main goal of our work is to resolve localization errors inherent in the existing bounding box labels. In other words, we consider a generic object detection scenario, where missing bounding boxes or incorrect class labels of the ground truth are uncommon. Instead, we assume that the bounding box labels may not be optimal due to the annotators’ mistakes or dilemma situations such as how to rotate a bounding box for a circular object on overhead imagery during the annotation process. In particular, we study a data augmentation strategy based on bounding box transformation for more robust localization.
@@ II-C Advanced options for translation and scaling
As demonstrated in Algorithm 1, the bounding box may be translated by different values along both the x-axis and y-axis. Besides, the aspect ratio may not be maintained for scaling operations. This phenomenon would not be recommended for training on certain data scenarios. For instance, it would not be desirable to modify the aspect ratio of bounding boxes in an extreme situation where most objects are expected to have a square size. Hence, we alleviate this by providing additional parameters ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ in our algorithm. Here, ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ are boolean values to be set by users. One can effortlessly modify NBBOX with ${\text{bool}}_{s}$ and ${\text{bool}}_{t}$ as follows.
@@
II-D Scale-aware noise injection into bounding box
Tiny objects often occupy only a few pixels in an image, making their bounding boxes highly sensitive to even minor perturbations. For tiny objects, such transformations can lead to significant discrepancies between the transformed bounding box and the actual object, resulting in huge misalignment [24]. To fill this gap, we propose enhancing our algorithm by replacing line 2 in Algorithm 1 with:
$\begin{split}&{x}_{\text{c}},{y}_{\text{c}},w,h,\theta\leftarrow\mathbf{B}_{i};\\ &\text{if }w\leq\gamma\text{ or }h\leq\gamma:\text{continue},\\ \end{split}$ (3)
where $\gamma\in\mathbb{Z}$ is a hyper-parameter to tune. For example, $\gamma=0$ means applying random transformations to all bounding boxes, as in our original algorithm. In contrast, when $\gamma$ is a positive integer, it indicates that bounding boxes with a width or height smaller than $\gamma$ will remain unaltered. Hence, the proper value of $\gamma$ helps preserve the integrity of tiny objects, ensuring that their spatial properties are not compromised by transformations that may lead to misalignment or loss of information in remote sensing imagery. Note that our Python implementation with Numpy slightly deviates from our pseudo-code for efficient computation (See supp. material).
@@ III-A Experimental setup
Detection models. Traditionally, deep learning-based object detection models are typically separated into one-stage and two-stage methods. Nowadays, the distinction between anchor-based and anchor-free methodologies tends to be more emphasized when classifying detector models. This work adopts rotated Faster R-CNN [25] and FCOS [26], representative models in anchor-based two-stage and anchor-free one-stage detection studies, respectively. For all models, either ImageNet-pretrained ResNet-50 [27], a representative CNN model, or Swin-T [28], a representative Transformer model, is used for feature extraction. Additionally, FPN-1x [29] is incorporated as a neck to enhance multi-scale invariance.
@@
Training details. We implement our hypothesis using PyTorch and MMRotate [30]. All the models are trained for 25 epochs. For a fair comparison with other approaches, all the experiments are resumed with weights pre-trained for five epochs without any augmentation per each {detector, data} combination. SGD is used for optimization with momentum of weight 0.9, weight decay 1e-4, learning rate 1e-3, and batch size 8. Learning rate warm-up is used over the first 500 iterations, where the learning rate linearly increases from zero to 33% of 5e-3. L2-norm gradient clipping is utilized with a maximum norm of 35. We apply the pre-processing operations ‘RResize’, ‘RRandomFlip’, ‘Normalize’, and ‘Pad’ for training, while ‘MultiScaleFlipAug’ is also used for testing.
@@
Figure 3: Impact of scaling, rotation, and translation of bounding boxes on model performance: as we observe similar outcomes across different architectures and datasets, for brevity, we include only the result from Faster R-CNN with ResNet-50 trained on DIOR-R.
@@ III-C Impact of scaling, rotation, and translation in NBBOX
This subsection studies the model performance when exposed to varying levels of scaling, rotation, and translation of bounding boxes. One fact is that too much deformation will adversely affect model training. In other words, to mitigate the bad effects of NBBOX, we assume that (i) ${s}_{\text{min}}$ and ${s}_{\text{max}}$ should be close to 1; (ii) ${r}_{\text{min}}$ and ${r}_{\text{max}}$ should be close to 0; (iii) ${t}_{\text{min}}$ should be close to -1 and ${t}_{\text{max}}$ should be close to 1. This reduces the difficulty of parameter tuning. To validate this assumption, we conduct experiments on DIOR-R with Faster R-CNN and FCOS, both employing ResNet-50 as the backbone (See Fig. 3 and supp. material for details). Here we consider combinations of the following parameters: ${s}_{\text{min}}\in$0.5, 0.8, 0.9, 0.95, 0.99}, ${s}_{\text{max}}\in$1, 1.01, 1.05, 1.1, 1.2, 1.5}, $\{{r}_{\text{min}},{r}_{\text{max}}\}\in$±0.01, ±0.1, ±0.5, ±1, ±2, ±3, ±4, ±5, ±10}, and $\{{t}_{\text{min}},{t}_{\text{max}}\}\in$±1, ±2, ±3, ±4, ±5, ±10, ±20, ±30}. As a result, we find parameters that support our hypothesis as:
  • ${s}_{\text{min}}\in$0.95, 1}, ${s}_{\text{max}}\in$1, 1.01};
  • $\{{r}_{\text{min}},{r}_{\text{max}}\}$ = {±0.01} & ${t}_{\text{min}}\in$-1, -2}, ${t}_{\text{max}}\in$1, 2}.
Unless otherwise specified, we use ${s}_{\text{min}}$ = 0.99, ${s}_{\text{max}}$ = 1.01, $\{{r}_{\text{min}},{r}_{\text{max}}\}$ = {±0.01}, and $\{{t}_{\text{min}},{t}_{\text{max}}\}$ = {±1} for remaining.
Isotropic scaling and translation. When scaling with preserving aspect ratio (i.e. ${\text{bool}}_{s}$=True), all models show a slight increase in mAP compared to when it is not (i.e. ${\text{bool}}_{s}$=False) as shown in Table VI. Besides, in the case of translation noise, the models perform slightly better when isotropic translation is applied (i.e. ${\text{bool}}_{t}$=True) compared to when it is not (i.e. ${\text{bool}}_{t}$=False) as indicated in Table VI. Thus, experimental results on DIOR-R in Table VI suggest that isotropic scaling and translation may be important. In the remaining experiments, we use ${\text{bool}}_{s}$=True and ${\text{bool}}_{t}$=True.
TABLE I: Average mAP under isotropic/non-isotropic scaling and
@@
(With ResNet-50) Faster R-CNN FCOS
Scaling ${\text{bool}}_{s}$=True 60.12 (+0.05) 56.97 (+0.05)
${\text{bool}}_{s}$=False 60.07 (+0.00) 56.92 (+0.00)
Translation ${\text{bool}}_{t}$=True 60.37 (+1.08) 58.38 (+1.46)
${\text{bool}}_{t}$=False 59.29 (+0.00) 56.92 (+0.00)
@@
TABLE II: Impact of combining scaling, rotation, and translation of bounding boxes on DIOR-R in NBBOX.
(Bold: best, underlined: second-best, wavy-underlined: third-best in this paper)
(With ResNet-50) Faster R-CNN FCOS
Scaling (Best) 62.79 ± 0.05 60.00 ± 0.99
Rotation (Best) 62.87 ± 0.29 59.68 ± 1.03
Translation (Best) 62.78 ± 0.26 59.93 ± 0.91
Scaling + Rotation 62.92 ± 0.24 60.13 ± 0.97
Scaling + Translation 62.89 ± 0.30 60.02 ± 1.07
Translation + Rotation 62.80 ± 0.22 59.89 ± 1.06
Scaling + Rotation + Translation 63.18 ± 0.14 60.25 ± 1.00
TABLE III: Impact of injecting scale-aware noise on DIOR-R in NBBOX.
(With ResNet-50) $\gamma$ = 0 $\gamma$ = 8 $\gamma$ = 16 $\gamma$ = 32 $\gamma$ = 64
Faster R-CNN 63.18 ± 0.14 63.20 ± 0.20 63.21 ± 0.18 63.17 ± 0.17 62.93 ± 0.27
FCOS 60.25 ± 1.00 60.24 ± 1.18 60.27 ± 1.06 59.99 ± 1.30 59.74 ± 1.24
Combining scaling, rotation, and translation. There is a reasonable expectation that integrating scaling, rotation, and translation could synergistically improve the model’s robustness and generalization. Hence, we further conduct experiments to validate the effectiveness of such combinations. Precisely, we study different combinations of each transformation (with parameters that show the best): pairwise combinations and the complete combination. Table VII indicates the effectiveness of the combined transformations in enhancing the model’s performance. Notably, the improvements are particularly evident in the scenario when we simultaneously use scaling, rotation, and translation during training.
III-D Impact of scale-aware noise injection in NBBOX
Following III-C, this subsection particularly focuses on a parameter study of $\gamma$, where $\gamma$ is varied over the set {0, 8, 16, 32, 64} under the combination of scaling, rotation, and translation. For this, we perform experiments on DIOR-R using Faster R-CNN and FCOS, both configured with ResNet-50 as the backbone as in III-C. As demonstrated in Table III, both detectors achieve their highest accuracies at $\gamma=16$, suggesting that bounding box perturbation for tiny objects (less than 16 $\times$ 16 pixels) might not be recommended. Besides, as $\gamma$ increases beyond 16, the performance degrades, indicating that moderate noise injection can enhance the model’s robustness. Hence, we use $\gamma$ = 16 for the remaining experiments.
III-E Comparison with state-of-the-art methods
In this subsection, we compare our method with easy-to-get data augmentation methods based on image manipulation (image/patch erasing and mix) as illustrated in Table IV. For a DOTA dataset, we use the same hyper-parameters for DIOR-R. Table IV shows that our NBBOX achieves a competitive performance among the state-of-the-art augmentation methods. Its performance can be further enhanced with other methods, such as RandRotate. Notably, obtaining satisfactory results on the DOTA dataset using the same parameters derived from DIOR-R demonstrates that the proposed method requires minimal parameter tuning. In addition, our method is faster than others: training Faster R-CNN with ResNet-50 on DIOR-R takes an average of 9.63 minutes per epoch for our method, compared to 21.24 minutes for MosaicMix. Due to the page limit, more results and analysis are in the supp. material.